OpenStudy (gavin39):

Stephanie is student teaching at her local high school. She has to help her instructor write a test. She has to write a radical equation where the solution is extraneous. She also has to write a radical equation where the solution is non-extraneous. Help Stephanie write one radical equation where the solution is extraneous and another equation where the solution is non-extraneous. Using complete sentences, explain each step when solving to justify your examples

OpenStudy (gavin39):

Will gove medals!

OpenStudy (gavin39):

Anyone?

OpenStudy (anonymous):

Ok let's do this step by step :-)

OpenStudy (anonymous):

you have sqrt (2y '+ 7) + 4 = y, you take 4 to the other side and square, 2y + 7 = (y - 4)^2

OpenStudy (anonymous):

Do you get that?